x^4 - 3x^2 + 2 = 0
Answer by stanbon(48546)

Factor to get:
(x^2-1)(x^2-2)=0
(x-1)(x+1)(x-sqrt2)(x+sqrt2)=0
x=1 or x=-1 or x=sqrt2 or x=-sqrt2
Cheers,
Stan H.